Evaluate the expression for the given f(x,y)
f(2, -3) if f(x, y) = x^2+y^2

Answer by jim_thompson5910(35256)

Plug in x = 2 and y = -3. Then evaluate.


f(x, y) = x^2+y^2


f(2, y) = (2)^2+y^2 ... Plug in x = 2 (ie, wherever you see an x, replace it with 2)


f(2, -3) = (2)^2+(-3)^2 ... Plug in y = -3 (ie, wherever you see a y, replace it with -3)


f(2, -3) = 4+9


f(2, -3) = 13